A 1.5-m wide pathway is constructed around the outside of a garden bed 15m by 18m. What is the area of the pathway.(round to the nearest square metre)
Answer by nerdybill(7384)

A 1.5-m wide pathway is constructed around the outside of a garden bed 15m by 18m. What is the area of the pathway.(round to the nearest square metre)
.
Garden area = 15*18 = 270 square meters
Total area = (15+2*1.5)(18+2*1.5) = = (15+3)(18+3) = (18)(21) = 378 square meters
.
area of pathway = "total area" - "garden area"
area of pathway = 378 - 270 = 108 square meters